# Homological shifts of a complementary edge ideal

### v1

@@ 1. Preliminary
Let $I\subseteq R$ be a monomial ideal with minimal multigraded free $S$-resolution
@@
where $F_{i}=\bigoplus\limits_{j=1}^{\beta_{i}(I)}S(-{\mathbf{a}}_{i,j})$. The vectors ${\mathbf{a}}_{i,j}\in{\mathbb{Z}}^{n}$ are called the $i$th multigraded shifts of the resolution $\mathbb{F}$. The monomial ideal
$\mbox{HS}_{i}(I)=\ ({\mathbf{x}}^{{\mathbf{a}}_{i,j}}:\ j=1,\dots,\beta_{i}(I))$
is called the $i$th homological shift ideal of $I$.
Note that $\mbox{HS}_{0}(I)=I$ and $\mbox{HS}_{i}(I)=(0)$ for $i<0$ or $i>\operatorname{pd}(I)$.
Let $I$ be a monomial ideal, and let $\mathcal{G}(I)$ be its unique minimal set of monomial generators. We say that $I$ has linear quotients if there exists an order $u_{1},\ldots,u_{m}$ of $\mathcal{G}(I)$ such that the ideal $(u_{1},\ldots,u_{i-1}):(u_{i})$ is generated by variables, for all $i=2,\ldots,m$. By [29, Lemma 2.1], we can always assume that $\deg(u_{1})\leq\dots\leq\deg(u_{m})$. When $I$ has linear quotients, we set $\operatorname{set}(u_{1})=\emptyset$ and
$\operatorname{set}(u_{i})\ =\ \{\ell\ :\ x_{\ell}\in(u_{1},\dots,u_{i-1}):(u_{i})\},\quad\textup{for}\ i=2,\dots,m.$
By [27, Lemma 1.5], if $I\subset S$ has linear quotients, then
(1) $\operatorname{HS}_{i}(I)\ =\ ({\bf x}_{F}u:\ u\in\mathcal{G}(I),\ F\subset\operatorname{set}(u),\ |F|=i),$
where ${\bf x}_{F}=\prod_{j\in F}x_{j}$ if $F\neq\emptyset$, and ${\bf x}_{\emptyset}=1$ otherwise. From this, we can conclude that the projection dimension of $I$ is
(2) $\mathrm{pd}(I)\ =\max\{|\operatorname{set}(u)|:\ u\in\mathcal{G}(I)\}.$
Let $G$ be a finite simple graph with vertex set $V(G)=\{1,\ldots,n\}$ and edge set $E(G)$. The complementary edge ideal $I_{c}(G)$ of $G$ is defined as the following monomial ideals
@@
The following Lemma can be proved in the same way as [23, Lemma 10.3.5].
@@
Let $I,J\subset S$ be monomial ideals with $d$-linear resolution and suppose that $J\subset I$. Then, for all ${\mathbf{a}}=(a_{1},\ldots,a_{n})\in\mathbb{N}^{n}$ and all $i\geq 0$, we have
@@
This result has two consequences. For a graph $G$, a subgraph $H$ of $G$ is called a spanning subgraph if $V(H)=V(G)$.
@@
Let $G$ be a connected graph and $H$ be a connected subgraph of $G$. Let $u=\prod_{i\in V(G)\setminus V(H)}x_{i}$. Then $u^{s}\cdot\operatorname{HS}_{i}(I_{c}(H)^{s})\subseteq\operatorname{HS}_{i}(I_{c}(G)^{s})$ for all $i,s\geq 1$. In particular, $\mathrm{pd}(I_{c}(H)^{s})\leq\mathrm{pd}(I_{c}(G)^{s})$.
@@
Let $G$ be a connected graph. Then $\beta_{i}(I_{c}(G)^{s})\leq\beta_{i}(I_{c}(G)^{s+1})$ for all $s\geq 1$. In particular, the projective dimension function $\mathrm{pd}(I_{c}(G)^{s})$ of $I_{c}(G)$ is a monotonically increasing function.
@@
Fix a minimal monomial generator $u$ of $I_{c}(G)$. Then $uI_{c}(G)^{s}\subseteq I_{c}(G)^{s+1}$ and both have (n2)(s+1)(n-2)(s+1)-linear resolution. It follows that $\beta_{i}(I_{c}(G)^{s})=\beta_{i}(uI_{c}(G)^{s})\leq\beta_{i}(I_{c}(G)^{s+1})$ by Lemma 1.2.    ∎
@@ 2. A Theorem on the homological shift ideals of monomial ideals with linear resolutions
Consider the short exact sequence of multi-graded $S$-modules:
$0\rightarrow\mathfrak{m}I\rightarrow I\rightarrow T\rightarrow 0,$
where $T=\bigoplus\limits_{X^{\alpha}\in\mathcal{G}(I)}\mathbb{K}[-\alpha]$.
Fix an integer $i\in[n]$ and a multi-degree $\beta\in\mathbb{N}^{n}$ with $|\beta|=i+d$. Since $I$ has a $d$-linear resolution, its $k$-th syzygy module is concentrated in degree $d+k$ for all $k\geq 0$, so $\mathrm{Tor}_{i-1}(I,\mathbb{K})_{\beta}=0$. For $\mathfrak{m}I$, note that multiplying an ideal with a $d$-linear resolution by $\mathfrak{m}$ yields an ideal with a $(d+1)$-linear resolution, so $\mathrm{Tor}_{i}(\mathfrak{m}I,\mathbb{K})_{\beta}=0$.
Applying the $\mathbb{K}\otimes_{S}$- functor to the short exact sequence above, we obtain the following long exact sequence in multi-graded $\mathbb{K}$-vector spaces (truncated at the relevant degrees):
$\rightarrow\mathrm{Tor}_{i}(\mathfrak{m}I,\mathbb{K})_{\beta}\rightarrow\mathrm{Tor}_{i}(I,\mathbb{K})_{\beta}\rightarrow\mathrm{Tor}_{i}(T,\mathbb{K})_{\beta}\rightarrow\\ \mathrm{Tor}_{i-1}(\mathfrak{m}I,\mathbb{K})_{\beta}\rightarrow\mathrm{Tor}_{i-1}(I,\mathbb{K})_{\beta}\rightarrow.$
@@
On the other hand, the Koszul complex of $x_{1},\ldots,x_{n}$ provides a minimal free resolution of $\mathbb{K}$. It follows that for each $X^{\alpha}\in\mathcal{G}(I)$, $\mathrm{Tor}_{i}(\mathbb{K}[-\alpha],\mathbb{K})_{\beta}\neq 0$ if and only if $\beta=\alpha+\gamma$ for some $\gamma\in\mathbb{N}^{n}$ with $X^{\gamma}\in\mathcal{G}(\mathfrak{m}^{[i]})$. Since $T=\bigoplus_{X^{\alpha}\in\mathcal{G}(I)}\mathbb{K}[-\alpha]$, $\mathrm{HS}_{i}(T)$ is generated by all monomials $X^{\alpha+\gamma}=X^{\alpha}X^{\gamma}$ where $X^{\alpha}\in\mathcal{G}(I)$ and $X^{\gamma}\in\mathcal{G}(\mathfrak{m}^{[i]})$, which is exactly $\mathfrak{m}^{[i]}I$.
Combining the two equalities, we get $\mathrm{HS}_{i}(I)+\mathrm{HS}_{i-1}(\mathfrak{m}I)=\mathfrak{m}^{[i]}I$.    ∎
This result indicates that every minimal generator of $\mathrm{HS}_{i}(I)$ has the form ${\bf x}_{F}\cdot u$, where $u\in\mathcal{G}(I)$ and $F$ is a subset of cardinality $i$ in $[n]$. This is consistent with the case when $I$ has linear quotients. In [14], the authors expected that if $I$ has linear resolutions then $\mathrm{HS}_{1}(I)$ would have linear quotients (see [14, the third paragraph, p134]). This result lends support to this expectation.
@@
Note that $\mathrm{HS}_{n}(I)=0$. Specializing $i$ to $n$ in Theorem 2.1 yields the desired formula.    ∎
3. Projective dimensions functions
In this section, we determine the projective dimension of powers of complementary edge ideals for several classes of connected graphs, including trees and cycles.
@@ 3. Projective dimensions functions
([1, Definition 6.2]) Two vertices $j$ and $k$ ($j$ may be equal to $k$) of $G$ are said to be even-connected with respect to the $s$-fold product $e_{1}\cdots e_{s}$ if there exists a walk $j_{1}--j_{2}--\cdots--j_{2\ell+2}$ ($\ell\geq 2$) in $G$ such that
@@
Let $G$ be a connected graph with vertex set $[n]$. As noted in [15, Remark 3.3], if the vertices of $G$ are labeled with integers from $[n]$ such that for all $i\in[n-1]$, the induced subgraph of $G$ on $V(G)\setminus[i]$ (i.e., on vertex set $\{i+1,\ldots,n\}$) is connected, then $I_{c}(G)^{k}$ has linear quotients for all $k\geq 1$ with respect to the lexicographical order on the minimal generators of $I_{c}(G)^{k}$ induced by $x_{1}>x_{2}>\cdots>x_{n}$. For the rest of this section, any order involved refers to this lexicographical order.
@@
Let $\beta\in\mathcal{G}(I_{c}(G)^{s})$, and write $\beta=\frac{\alpha^{s}}{u_{1}\cdots u_{s}}$, where each $u_{i}\in\mathcal{G}(I(G))$, and $I(G)$ is the edge ideal of $G$. Then
@@
For any monomials $U_{1},\ldots,U_{k},U$, one has $(\frac{\alpha^{s}}{U_{1}},\cdots,\frac{\alpha^{s}}{U_{k}}):\frac{\alpha^{s}}{U}=(U:U_{1})+\cdots+(U:U_{k})$. The first assertion follows from this observation. Next, we prove the second assertion.
Let $u=u_{1}\cdots u_{s}$. If $i\in\mathrm{set}(\beta)$, then there is some $\gamma>\beta$ such that $\gamma=\alpha^{s}/v$ and $\gamma:\beta=u:v=(x_{i})$, where $v\in\mathcal{G}(I(G)^{s})$. Note that $u=x_{i}x_{j}u_{2}\cdots u_{s}$, we can deduce that $v=x_{k}x_{j}u_{2}\cdots u_{s}$ for some $k\in[n]$ and that $u>v$, which implies that $i<k$. It is clear that $x_{k}x_{j}\in(I(G)^{s}:u_{2}\cdots u_{s})$, so by Theorem 3.2, either $\{j,k\}\in E(G)$ or $j$ is even-connected to $k$ with respect to the (s1)(s-1)-fold product $u_{2}\cdots u_{s}$.
Conversely, if there is $k\in[n]$ such that $k>i$, and either $\{j,k\}\in E(G)$ or $j$ is even-connected to $k$ with respect to the (s1)(s-1)-fold product $u_{2}\cdots u_{s}$, then $x_{k}x_{j}\in(I(G)^{s}:u_{2}\cdots u_{s})$. Let $v=x_{k}x_{j}u_{2}\cdots u_{s}$ and $\gamma=\alpha^{s}/v$. Then $\gamma>\beta$ and $\gamma\in\mathcal{G}(I_{c}(G)^{s})$. Clearly, $\gamma:\beta=(x_{i})$, so $i\in\mathrm{set}(\beta)$.    ∎
By [17, Propositions 4.3, 4.5], if $G$ is a connected graph, then the maximal value of $\mathrm{pd}(I_{c}(G)^{s})$ for $s\geq 1$ is n1n-1 if $G$ is non-bipartite, and n2n-2 if $G$ is bipartite. We first determine the projective dimension function of the complementary edge ideal of a cycle.
@@
(1) Suppose the vertices of $G$ are labeled as $1-2-3-\cdots-(2m-1)-(2m)-1$. Fix an $s\in[m-1]$. For any $\beta=\frac{\alpha^{s}}{u_{1}\cdots u_{s}}\in\mathcal{G}(I_{c}(G)^{s})$ where each $u_{i}\in\mathcal{G}(I(G))$, Lemma 3.3 implies $\mathrm{set}(\beta)\subseteq\{i\mid x_{i}\mbox{ divides }u_{1}u_{2}\cdots u_{s}\}$. Thus $|\mathrm{set}(\beta)|\leq 2s$, so $\mathrm{pd}(I_{c}(G)^{s})=\max\{|\mathrm{set}(\alpha)|:\alpha\in\mathcal{G}(I_{c}(G)^{s})\}\leq 2s$.
Conversely, we choose $u_{i}=x_{2i-1}x_{2i}$ for any $i\in[s]$, and set $\beta=\frac{\alpha^{s}}{u_{1}\cdots u_{s}}$. Then $\beta\in\mathcal{G}(I_{c}(G)^{s})$. By Lemma 3.3, $(2i-1)\in\mathrm{set}(\beta)$ for any $i\in[s]$. Note that $2m$ is even-connected to 2i12i-1 with respect to $u_{1}u_{2}\cdots u_{2i-3}$ via the walk:
$(2m)-1-2-3-4-\cdots-(2i-3)-(2i-2)-(2i-1).$
Given this fact and since $2m>2i$, we conclude $2i\in\mathrm{set}(\beta)$ by Lemma 3.3. Therefore, $\mathrm{set}(\beta)=[2s]$ and $\mathrm{pd}(I_{c}(G)^{s})=\max\{|\operatorname{set}(\alpha)|:\alpha\in\mathcal{G}(I_{c}(G)^{s})\}\geq 2s$.
Since $\mathrm{pd}(I_{c}(G)^{s})$ is increasing by Lemma 1.3, and $\mathrm{pd}(I_{c}(G)^{s})\leq 2m-2$ for all $s\geq 1$ by [17, Proposition 4.3], it follows that $\mathrm{pd}(I_{c}(G)^{s})=2m-2$ for all $s\geq m-1$.
The proof of (2) is similar to that of (1), so we omit the proof.    ∎
@@
Let $G$ be a tree with $n$ vertices. We fix a degree-one vertex of $G$ and label it $n$. There exists at least one labeling of the remaining vertices in $V(G)$ using the elements of [n1][n-1] such that $\text{dist}(i,n)\geq\text{dist}(j,n)$ for all $1\leq i<j\leq n-1$. Here, $\text{dist}(i,n)$ is the length of the shortest path between $i$ and $n$ (commonly referred to as “distance” in graph theory). In such a labeling, for each $i\in[n-1]$, there is a unique $k$ with $i<k\leq n$ such that $i$ is adjacent to $k$. This kind of labeling is not unique in general, and is unique precisely when $G$ is a path. In the rest of this section, all trees will be labeled in this manner.
@@
Let $G$ be a tree and $\beta\in\mathcal{G}(I_{c}(G)^{s})$. Write $\beta=\frac{\alpha^{s}}{u_{1}\cdots u_{s}}$, where each $u_{i}\in\mathcal{G}(I(G))$, Assume that $u_{1}=x_{i}x_{j}$ with $1\leq i<j\leq n$. Then

### v5

@@ 2.2. Core Concepts: Homological Shift Ideals and Algebras
Let $I\subset S$ be a monomial ideal with minimal multigraded free $S$-resolution
@@
where $F_{i}=\bigoplus\limits_{j=1}^{\beta_{i}(I)}S(-\mathbf{a}_{i,j})$. The vectors $\mathbf{a}_{i,j}\in\mathbb{Z}^{n}$ are called the $i$-th multigraded shifts of the resolution $\mathbb{F}$. The monomial ideal
$\mathrm{HS}_{i}(I):=\ (\mathbf{x}^{\mathbf{a}_{i,j}}:\ j=1,\dots,\beta_{i}(I))$
is called the $i$-th homological shift ideal of $I$.
Note that $\mathrm{HS}_{0}(I)=I$ and $\mathrm{HS}_{i}(I)=(0)$ for $i<0$ or $i>\operatorname{pd}(I)$. If a monomial ideal admits a $d$-linear resolution, then $\operatorname{HS}_{i}(I)$ is either 0 or generated in degree $i+d$ for all $i$.
Definition 2.2.
Let $I\subset S$ be a monomial ideal. We define its $i$-th homological shift algebra $\mathrm{HS}_{i}(\mathcal{R}(I))$ as
$\mathrm{HS}_{i}(\mathcal{R}(I)):=\bigoplus_{s\geq 0}\mathrm{HS}_{i}(I^{s}).$
In general, $\mathrm{HS}_{i}(\mathcal{R}(I))$ is merely an abelian group. However, if $I^{s}$ admits a linear resolution for all $s\geq 1$, then $\mathrm{HS}_{i}(\mathcal{R}(I))$ is a finitely generated graded $\mathcal{R}(I)$-module.
2.3. Linear Quotients and Related Properties
Linear quotients are a key tool for characterizing the structure of homological shift ideals. We recall its definition and a fundamental formula below.
Let $I$ be a monomial ideal, and let $\mathcal{G}(I)$ denote its unique minimal set of monomial generators. We say that $I$ has linear quotients if there exists an ordering $u_{1},\ldots,u_{m}$ of $\mathcal{G}(I)$ such that the colon ideal $(u_{1},\ldots,u_{i-1}):(u_{i})$ is generated by variables for all $i=2,\ldots,m$. By [28, Lemma 2.1], we may assume without loss of generality that $\deg(u_{1})\leq\cdots\leq\deg(u_{m})$.
When $I$ admits linear quotients, we define $\operatorname{set}(u_{1})=\emptyset$ and
$\operatorname{set}(u_{i})=\left\{\ell\ :\ x_{\ell}\in(u_{1},\ldots,u_{i-1}):(u_{i})\right\},\quad\text{for }i=2,\ldots,m.$
By [26, Lemma 1.5], it follows that
(1) $\operatorname{HS}_{i}(I)=\left(\mathbf{x}_{F}u:\ u\in\mathcal{G}(I),\ F\subseteq\operatorname{set}(u),\ |F|=i\right),$
where $\mathbf{x}_{F}=\prod_{j\in F}x_{j}$ if $F\neq\emptyset$, and $\mathbf{x}_{\emptyset}=1$ otherwise. From this formula, we deduce that the projective dimension of $I$ satisfies
(2) $\mathrm{pd}(I)=\max\left\{|\operatorname{set}(u)|:\ u\in\mathcal{G}(I)\right\}.$
2.4. Complementary edge ideal and FM condition
Let $G$ be a finite simple graph with vertex set $V(G)=\{1,\ldots,n\}$ and edge set $E(G)$. The complementary edge ideal $I_{c}(G)$ of $G$ is defined as the monomial ideal
@@ 2.4. Complementary edge ideal and FM condition
A key characterization of $I_{c}(G)$ is as follows: $I_{c}(G)$ admits a linear resolution if and only if $G$ is connected, if and only if $I_{c}(G)^{s}$ admits linear quotients for all $s\geq 1$. In [12], the ordering of minimal generators of $I_{c}(G)^{s}$ that yields linear quotients was given as follows.
For a simple graph $G=(V,E)$ and a vertex subset $U\subseteq V$, the induced subgraph of $G$ on $U$, denoted $G[U]$, is the graph with vertex set $U$ and edge set $\{\{u,v\}\in E\mid u,v\in U\}$.
Definition 2.3.
Let $G$ be a connected graph with $n$ vertices. A labeling of $V(G)$ with $[n]$ is said to satisfy the FM condition if, for all $i\in[n-1]$, the induced subgraph $G[V(G)\setminus[i]]=G[\{i+1,\dots,n\}]$ is connected. This condition is introduced in the work of Ficarra and Moradi [12, Section 3]; we adopt the name FM condition here for brevity, following the initials of the authors.
As noted in [12, Remark 3.3], if $V(G)$ admits a labeling by $[n]$ that satisfies the FM condition, then $I_{c}(G)^{s}$ admits linear quotients for all $s\geq 1$ with respect to the lexicographical order on the minimal generators of $I_{c}(G)^{s}$ induced by $x_{1}>x_{2}>\cdots>x_{n}$. Throughout this paper, any ordering of the minimal monomial generators of an ideal $I$ with respect to which $I$ admits linear quotients refers to this specific lexicographical order.
2.5. Key Lemmas
The following lemma can be proved in the same way as [21, Lemma 10.3.5].
@@ 2.5. Key Lemmas
Let $I,J\subset S$ be monomial ideals that admit $d$-linear resolutions and suppose that $J\subset I$. Then, for all $\mathbf{a}=(a_{1},\ldots,a_{n})\in\mathbb{N}^{n}$ and all $i\geq 0$, we have
@@
Lemma 2.4 yields the following two immediate corollaries, which are crucial to our subsequent arguments.
@@
Let $G$ be a connected graph and $H$ be a connected subgraph of $G$. Let $u:=\prod_{i\in V(G)\setminus V(H)}x_{i}$. Then $u^{s}\cdot\operatorname{HS}_{i}(I_{c}(H)^{s})\subseteq\operatorname{HS}_{i}(I_{c}(G)^{s})$ for all $i,s\geq 1$. In particular, $\mathrm{pd}(I_{c}(H)^{s})\leq\mathrm{pd}(I_{c}(G)^{s})$.
Proof.
Note that $u^{s}I_{c}(H)^{s}\subseteq I_{c}(G)^{s}$. By the characterization of complementary edge ideals, both $u^{s}I_{c}(H)^{s}$ and $I_{c}(G)^{s}$ admit linear resolutions. The result thus follows immediately from Lemma 2.4.
@@
Let $G$ be a connected graph. Then $\beta_{i}(I_{c}(G)^{s})\leq\beta_{i}(I_{c}(G)^{s+1})$ for all $s\geq 1$. In particular, the projective dimension function $\mathrm{pd}(I_{c}(G)^{s})$ of $I_{c}(G)$ is a non-decreasing function of $s$.
@@
Fix a minimal monomial generator $u$ of $I_{c}(G)$. Then $uI_{c}(G)^{s}\subseteq I_{c}(G)^{s+1}$, and both ideals admit (n2)(s+1)(n-2)(s+1)-linear resolutions. It follows that $\beta_{i}(I_{c}(G)^{s})=\beta_{i}(uI_{c}(G)^{s})\leq\beta_{i}(I_{c}(G)^{s+1})$ by Lemma 2.4.
@@ 3. A Theorem on the homological shift ideals of monomial ideals with linear resolutions
Let $T:=\bigoplus\limits_{{\bf x}^{\alpha}\in\mathcal{G}(I)}\mathbb{K}[-\alpha]$, then we obtain the short exact sequence of multigraded $S$-modules:
$0\rightarrow\mathfrak{m}I\rightarrow I\rightarrow T\rightarrow 0.$
Fix an integer $i\in[n]$ and a multidegree $\beta\in\mathbb{N}^{n}$ with $|\beta|=i+d$. Since $I$ has a $d$-linear resolution, its $k$-th syzygy module is concentrated in degree $d+k$ for all $k\geq 0$, so $\mathrm{Tor}_{i-1}(I,\mathbb{K})_{\beta}=0$. For $\mathfrak{m}I$, note that multiplying an ideal with a $d$-linear resolution by $\mathfrak{m}$ yields an ideal with a $(d+1)$-linear resolution, so $\mathrm{Tor}_{i}(\mathfrak{m}I,\mathbb{K})_{\beta}=0$.
Applying the $\mathbb{K}\otimes_{S}$- functor to the short exact sequence above, we obtain the following long exact sequence of multigraded $\mathbb{K}$-vector spaces (truncated at the relevant degrees):
$\rightarrow\mathrm{Tor}_{i}(\mathfrak{m}I,\mathbb{K})_{\beta}\rightarrow\mathrm{Tor}_{i}(I,\mathbb{K})_{\beta}\rightarrow\mathrm{Tor}_{i}(T,\mathbb{K})_{\beta}\rightarrow$
$\quad\rightarrow\mathrm{Tor}_{i-1}(\mathfrak{m}I,\mathbb{K})_{\beta}\rightarrow\mathrm{Tor}_{i-1}(I,\mathbb{K})_{\beta}\rightarrow.$
@@
On the other hand, the Koszul complex of $x_{1},\ldots,x_{n}$ provides a minimal free resolution of $\mathbb{K}$. It follows that for each ${\bf x}^{\alpha}\in\mathcal{G}(I)$, $\mathrm{Tor}_{i}(\mathbb{K}[-\alpha],\mathbb{K})_{\beta}\neq 0$ if and only if $\beta=\alpha+\gamma$ for some $\gamma\in\mathbb{N}^{n}$ with ${\bf x}^{\gamma}\in\mathcal{G}(\mathfrak{m}^{[i]})$. Since $T=\bigoplus_{{\bf x}^{\alpha}\in\mathcal{G}(I)}\mathbb{K}[-\alpha]$, $\mathrm{HS}_{i}(T)$ is generated by all monomials ${\bf x}^{\alpha+\gamma}={\bf x}^{\alpha}{\bf x}^{\gamma}$ where ${\bf x}^{\alpha}\in\mathcal{G}(I)$ and ${\bf x}^{\gamma}\in\mathcal{G}(\mathfrak{m}^{[i]})$, which is exactly $\mathfrak{m}^{[i]}I$.
Combining the two equalities, we get $\mathrm{HS}_{i}(I)+\mathrm{HS}_{i-1}(\mathfrak{m}I)=\mathfrak{m}^{[i]}I$.
This result indicates that every minimal generator of $\mathrm{HS}_{i}(I)$ has the form ${\bf x}_{F}\cdot u$, where $u\in\mathcal{G}(I)$ and $F$ is a subset of cardinality $i$ in $[n]$. This is consistent with the case when $I$ has linear quotients. In [11], the authors expected that if $I$ has linear resolutions then $\mathrm{HS}_{1}(I)$ would have linear quotients (see [11, the third paragraph, p134]). This result lends support to this expectation.
@@
Note that $\mathrm{HS}_{n}(I)=0$. Specializing $i$ to $n$ in Theorem 3.1 yields the desired formula.
4. Projective dimension functions
In this section, we investigate the projective dimension of powers of complementary edge ideals for several classes of connected graphs, including trees and cycles.
@@ 4. Projective dimension functions
Recall that a walk in a graph $G$ is a finite sequence of vertices $v_{0},v_{1},\dots,v_{k}$ such that for every $i=1,2,\dots,k$, there is an edge between $v_{i-1}$ and $v_{i}$ in $G$; such a walk is compactly denoted by $v_{0}-v_{1}-\cdots-v_{k}$. The length of this walk is defined to be the number of edges it contains, i.e., $k$. A walk is called a closed walk if its starting vertex coincides with its ending vertex, i.e., $v_{0}=v_{k}$. Vertices and edges can be repeated in a walk. By contrast, a path in a graph $G$ is a walk where all vertices are distinct.
@@
([1, Definition 6.2]) Two vertices $j$ and $k$ ($j$ may be equal to $k$) of $G$ are said to be even-connected with respect to the $s$-fold product $e_{1}\cdots e_{s}$ of edges if there exists a walk $j_{1}-j_{2}-\cdots-j_{2\ell+2}$ ($\ell\geq 2$) in $G$ such that
@@
Let $G$ be a connected graph whose vertices is labeled in a way satisfying the FM condition. Let $\beta\in\mathcal{G}(I_{c}(G)^{s})$, and write $\beta=\frac{\alpha^{s}}{u_{1}\cdots u_{s}}$, where each $u_{i}\in\mathcal{G}(I(G))$. Then
@@
For any monomials $U_{1},\ldots,U_{k},U$ that divides $\alpha^{s}$, one has $(\frac{\alpha^{s}}{U_{1}},\cdots,\frac{\alpha^{s}}{U_{k}}):\frac{\alpha^{s}}{U}=(U:U_{1})+\cdots+(U:U_{k})$. The first assertion follows from this observation. Next, we prove the second assertion.
Let $u:=u_{1}\cdots u_{s}$. If $i\in\mathrm{set}(\beta)$, then there exists some $\gamma>\beta$ such that $\gamma=\alpha^{s}/v$ and $\gamma:\beta=u:v=(x_{i})$, where $v\in\mathcal{G}(I(G)^{s})$.
Since $u=x_{i}x_{j}u_{2}\cdots u_{s}$, we deduce $v=x_{k}x_{j}u_{2}\cdots u_{s}$ for some $k\in[n]$. As $\gamma>\beta$, it follows that $u>v$, which in turn implies $i<k$. It follows directly that $x_{k}x_{j}\in(I(G)^{s}:u_{2}\cdots u_{s})$. By Theorem 4.2, either $\{j,k\}\in E(G)$ or $j$ is even-connected to $k$ with respect to the (s1)(s-1)-fold product $u_{2}\cdots u_{s}$.
Conversely, suppose there exists $k\in[n]$ such that $k>i$, and either $\{j,k\}\in E(G)$ or $j$ is even-connected to $k$ with respect to the (s1)(s-1)-fold product $u_{2}\cdots u_{s}$. Then $x_{k}x_{j}\in(I(G)^{s}:u_{2}\cdots u_{s})$. Let $v:=x_{k}x_{j}u_{2}\cdots u_{s}$ and $\gamma:=\alpha^{s}/v$. It follows that $\gamma>\beta$ and $\gamma\in\mathcal{G}(I_{c}(G)^{s})$. Moreover, $\gamma:\beta=(x_{i})$, which implies $i\in\mathrm{set}(\beta)$.
By [14, Propositions 4.3, 4.5], if $G$ is a connected graph, then the maximal value of $\mathrm{pd}(I_{c}(G)^{s})$ for $s\geq 1$ is n1n-1 if $G$ is non-bipartite, and n2n-2 if $G$ is bipartite. We first determine the projective dimension function of the complementary edge ideal of a cycle.
@@
(1) Suppose the vertices of $G$ are labeled consecutively as $1,2,3,\dots,2m-1,2m$ in a cyclic order. It is clear that this labeling satisfies the FM condition (see Definition 2.3). Fix an integer $s\in[m-1]$. Then, for any $\beta=\frac{\alpha^{s}}{u_{1}u_{2}\cdots u_{s}}\in\mathcal{G}(I_{c}(G)^{s})$ where each $u_{i}\in\mathcal{G}(I(G))$, Lemma 4.3 ensures that $\mathrm{set}(\beta)\subseteq\left\{i\mid x_{i}\text{ divides }u_{1}u_{2}\cdots u_{s}\right\}$. It follows immediately that $|\mathrm{set}(\beta)|\leq 2s$. Consequently, by formula (2), we have
$\mathrm{pd}(I_{c}(G)^{s})=\max\left\{|\mathrm{set}(\beta)|:\beta\in\mathcal{G}(I_{c}(G)^{s})\right\}\leq 2s.$
Conversely, let $\mathbf{e}_{i}:=x_{2i-1}x_{2i}$ for each $i\in[s]$, and define $\gamma:=\frac{\alpha^{s}}{\mathbf{e}_{1}\mathbf{e}_{2}\cdots\mathbf{e}_{s}}$. Then $\gamma\in\mathcal{G}(I_{c}(G)^{s})$. An application of Lemma 4.3 gives $2i-1\in\mathrm{set}(\gamma)$ for all $i\in[s]$. Next, note that vertex $2m$ is even-connected to vertex 2i12i-1 with respect to $\mathbf{e}_{1}\mathbf{e}_{2}\cdots\mathbf{e}_{2i-3}$, via the walk:
$2m-1-2-3-4-\cdots-(2i-3)-(2i-2)-(2i-1).$
By virtue of this adjacency walk and the inequality $2m>2i$, a further appeal to Lemma 4.3 yields $2i\in\mathrm{set}(\gamma)$. We thus have $\mathrm{set}(\gamma)=[2s]$, from which it follows that $\mathrm{pd}(I_{c}(G)^{s})=\max\left\{|\mathrm{set}(\beta)|:\beta\in\mathcal{G}(I_{c}(G)^{s})\right\}\geq 2s.$
Combining the upper and lower bounds, we establish the equality $\mathrm{pd}(I_{c}(G)^{s})=2s$ for all integers $s\in[m-1]$.
Since $\mathrm{pd}(I_{c}(G)^{s})$ is a non-decreasing function of $s$ by Lemma 2.5, while $\mathrm{pd}(I_{c}(G)^{s})\leq 2m-2$ for all $s\geq 1$ by [14, Proposition 4.3], it follows that $\mathrm{pd}(I_{c}(G)^{s})=2m-2$ for all $s\geq m-1$.
The proof of (2) is similar to that of (1), so we omit the proof.
Next, we turn our attention to trees. To begin with, we assign $V(G)$ a labeling for an arbitrary tree $G$ that satisfies the FM condition as specified in Definition 2.3.
@@
Let $G$ be a tree with $n$ vertices. We fix a vertex of degree one as the root and label it $n$. To label the remaining n1n-1 vertices with [n1][n-1], we use a distance-based hierarchical rule as follows:
For any vertex $u\in V(G)$, let $\text{dist}(u,n)$ denote the length of the unique shortest path from $u$ to $n$. Group $V(G)\setminus\{n\}$ into layers $L_{d}=\{u\mid\text{dist}(u,n)=d\}$ for $d\geq 1$. Denote $m:=\max\{d\mid L_{d}\neq\emptyset\}$. Assign labels in [n1][n-1] from the farthest layer to the closest, that is, from $L_{m}$ down to $L_{1}$: assign the smallest consecutive labels to $L_{m}$, the next set to $L_{m-1}$, and so on. The labeling order within the same layer is arbitrary.
It is easy to verify that such a labeling satisfies the FM condition. Furthermore, a key property holds: For each $j\in[n-1]$, there exists a unique $j<k\leq n$ such that $j$ and $k$ are adjacent. We call $k$ the “parent” of $j$ relative to root $n$.
This labeling is generally non-unique since labels within a layer can be permuted, but is unique if and only if $G$ is a path. In the remainder of this paper, all trees are labeled in this manner.
@@
Let $G$ be a tree and let $i-y_{1}-y_{2}-\cdots-y_{p}-k$ be a walk in $G$ with $i<k$. If $y_{1}<i$, then there exists $2\leq q\leq p$ such that $y_{q}=i$.
Proof.
First, define a function $\phi:[n-1]\rightarrow[n]$ where $\phi(j)$ denotes the unique parent of $j$ (as defined in Remark 4.5): for each $j\in[n-1]$, $\phi(j)$ is the unique vertex $k>j$ that is adjacent to $j$. A fundamental observation about $\phi$ is that for any walk in $G$ starting at a vertex $j\in[n-1]$ and ending at a vertex $\ell>j$, the parent $\phi(j)$ must lie on this walk. This follows from the tree structure: the unique path from $j$ to $\ell$ (and hence any walk from $j$ to $\ell$, which reduces to this path) must pass through $\phi(j)$, the unique neighbor of $j$ with a larger label.
Applying this observation to the walk $y_{1}-y_{2}-\cdots-y_{p}-k$ and noting that $\phi(y_{1})=i$, the result follows.
Lemma 4.7.
Let $G$ be a tree and $\beta\in\mathcal{G}(I_{c}(G)^{s})$. Write $\beta=\frac{\alpha^{s}}{u_{1}\cdots u_{s}}$, where each $u_{i}\in\mathcal{G}(I(G))$; assume that $u_{1}=x_{i}x_{j}$ with $1\leq i<j\leq n$. Then